### v1

# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Intermediate Feature Distance

Abstract
Deep Neural Networks (DNNs) are vulnerable to adversarial attacks, which pose security challenges to hyperspectral image (HSI) classification technologies based on DNNs. In the domain of natural images, numerous transfer-based adversarial attack methods have been studied, among which input transformation is a common strategy to mitigate the overfitting of adversarial examples to surrogate models by applying diverse transformations to the images. Feature-space adversarial attacks, which modulate intermediate-layer representations of images, also demonstrate strong transferability. However, HSIs differ from natural images due to their high-dimensional and rich spectral information. Current research on HSI adversarial examples remains limited and faces challenges in fully utilizing the structural and feature information of images. To address these challenges, this paper proposes a novel method to enhance the transferability of the adversarial examples for HSI classification models. First, while keeping the image structure unchanged, the proposed method randomly divides the image into blocks in both spatial and spectral dimensions. Then, various transformations are applied on a block by block basis to increase input diversity and mitigate overfitting. Second, a feature distancing loss targeting intermediate layers is designed, which measures the distance between the amplified features of the original examples and the features of the adversarial examples as the primary loss, while the output layer prediction serves as the auxiliary loss. This guides the perturbation to disrupt the features of the true class in adversarial examples, effectively enhancing transferability. Extensive experiments demonstrate that the adversarial examples generated by the proposed method achieve effective transferability to black-box models on two public HSI datasets. Furthermore, the method maintains robust attack performance even under defense strategies.
Index Terms:
Hyperspectral image classification, Adversarial examples, Transferability, Feature distancing
I Introduction
While obtaining geometric features of object surfaces through two-dimensional spatial imaging technology, hyperspectral images (HSIs) employ spectral demodulation techniques to interpret the total radiant energy of each pixel as a continuous spectrum across different bands [1]. Compared with natural images that have only three RGB channels, HSIs build a three-dimensional data cube that integrates the spatial-spectral information of ground targets, enabling the capture of more subtle differences between them. Due to this advantage, HSIs have important application value in meteorology [2], environmental protection [3], agriculture [4], and military fields [5].
As a fundamental task in HSI applications, HSI classification aims to identify the substances represented by hyperspectral pixels. In recent years, DNNs have become the dominant approach for HSI classification due to their powerful feature extraction capabilities [6, 7, 8, 9]. Researchers have developed various methods based on convolutional neural networks (CNNs), recurrent neural networks (RNNs), and other DNN models to extract more discriminative spatial and spectral features from HSIs [10, 11, 12]. For example, Wang et al. [10] introduced a dynamic super-pixel normalization (DSN)-based DNN for HSI classification; Mahmoodi et al. [11] developed an effective method capable of focusing on both spectral and spatial features. When considering special circumstances such as insufficient training samples, more methods have been presented based on the techniques like few-shot learning for HSI classification [13, 14, 15].
Despite demonstrating significant advantages in HSI classification tasks, DNNs’ vulnerability to adversarial attacks and the reliability risks posed by adversarial examples still require focused attention. Many studies [16, 17] show that DNNs are highly vulnerable to adversarial attacks, which cause machine learning models to produce erroneous predictions via imperceptible subtle perturbations added to input data. These crafted inputs are referred to as adversarial examples. Current research on adversarial attacks mainly focuses on white-box settings where adversaries possess the target model’s internal information (including architecture, parameters, and training data) and leverage such knowledge to design targeted attacks. In contrast, black-box attacks do not require such access and are thus more realistic. Among the black-box attacks [18, 19], transfer-based attacks [20, 21] which utilize the transferability of adversarial examples to perform attacks are more flexible, where transferability refers to the property that adversarial examples trained on a known substitute model are also effective against structurally different target models.
Researchers by now have paid much attention to adversarial examples in the natural image domain [22], with numerous methods proposed to enhance their transferability. These include classic approaches such as the FGSM method [23], and subsequent iterative methods [17, 24, 25]. However, adversarial examples generated using a substitute model often suffer from overfitting to the source model’s architecture and feature representations. To address this, various input transformation methods (e.g.,[26, 27, 28]) have been developed to improve transferability success rates. For example, the diverse input method (DI) [27] convert the input with a probability $p$, achieved by randomly resizing the image and padding zeros at its edges. Additionally, feature-space attack methods (e.g.,[21, 29, 30, 31]) aim to enhance transferability by modifying feature representations or altering the model’s attention, often guiding perturbation updates more purposefully through advanced loss functions. For example, Xu et al. [32] proposed a new black-box adversarial attack method called Mixup-Attack and its variant Mixcut-Attack, which finds common vulnerabilities between different networks by attacking the shallow features of a given substitute model.
However, due to the significant differences between HSIs and natural images, HSI classification needs to consider not only spatial but also spectral domain information, significantly increasing the complexity of generating perturbations. The data transformation-based attack methods proposed so far are generally more suitable for natural images. Performing transformations only in the spatial dimension fails to effectively utilize the redundant characteristics of HSI spectral data. Moreover, we argue that the structure of HSIs also contains a wealth of information, but previous methods (e.g., Shi et al. [33] approach utilizing superpixel segmentation and spectral clustering to enhance adversarial examples robustness and transferability) have not considered using this semantic information, leaving adversarial examples still suffer from the problem of overfitting to the substitute model. Furthermore, existing methods for generating adversarial examples usually rely on the classification layer information (e.g., using cross-entropy loss) to guide the generation of adversarial examples. However, the deep structure of neural networks leads to the problem of vanishing gradients during backpropagation [21], so relying solely on the distance between predicted labels as the optimization objective has significant limitations. Although methods like [32, 34] employ shallow-layer feature distances as losses, some models’ early layers may fail to learn truly salient features and semantic concepts [29], limiting the ability to guide perturbations in disrupting intra-image features and thus resulting in limited transferability.
Therefore, this study focuses on improving the transferability of adversarial examples in HSIs by enhancing input transformations and fully utilizing feature information. Wang et al. [35] observed in the natural image domain that more diverse image transformations can improve transferability while maintaining the structure unchanged. Given that HSIs typically contain abundant spectral bands (e.g.,the PaviaU dataset has 115 bands), we leverage this characteristic by dividing the input hyperspectral examples into blocks along both spatial and spectral dimensions and applying random transformations block by block to increase the diversity of the transformed images. Additionally, we perform attacks in the feature space to increase the distance between the original and adversarial examples. Inspired by the non-robust feature alignment method proposed by Zhu et al. [36] in the targeted attack approach, this paper designs a channel-weighted non-robust feature distancing loss (i.e., increasing inter-feature distances) from a non-targeted attack perspective. Initially, we calculate the weight of each channel to disrupt features that have a greater impact on model output. When implementing feature distance, we first enlarge the feature values of the original example, so that the feature values of the adversarial example become smaller than those of the original ones. This allows the perturbations to be updated in a direction that suppresses the original features within the adversarial examples. Finally, by integrating information from the model’s output layer, we construct a total loss function which forces the features of the adversarial example to deviate from the original distribution, thereby improving transferability.
Our contributions are summarized as follows:
  1. 1.
    We propose a novel attack framework to boost adversarial transferability for HSI classification by utilizing 3D structure-invariant transformations and intermediate feature distancing. While keeping the image structure unchanged, considering the numerous spectral bands of HSIs, we divided the image along both spatial and spectral dimensions. Subsequently, we applied random transformations to each block to enhance the diversity of the transformed images.
  2. 2.
    Previous methods mainly used classification layer information for loss calculation, resulting in insufficient utilization of image feature information. In this paper, we propose a new loss function that amplifies the features of the original examples and keeps them away from the features of adversarial examples. While calculating the information using channel weights, our method also integrates the classification information from the output layer to guide the exploration of more effective perturbations.
  3. 3.
    The transferability and defense experiments conducted on HSI datasets (PaviaU and HoustonU 2018) validate the effectiveness of the proposed method in generating adversarial examples with superior transferability and robustness against defenses.
The rest of this paper is organized as follows: Section 2 reviews related works relevant to this study. Section 3 provides a detailed explanation of the proposed adversarial attack methodology. Section 4 introduces the dataset information, experimental settings, and summarizes the experimental results. Section 5 concludes the study.
II Related Work
II-A Traditional White-box Attack Methods
Gradient-based methods are a common and effective category of adversarial attack techniques. The core idea is to maximize the model’s loss using its gradient information to generate adversarial examples. Goodfellow et al. [23] proposed the Fast Gradient Sign Method (FGSM), which generates powerful adversarial examples based on the linear properties of CNNs. Wang et al. [17] and Madry et al. [24] further extended FGSM by decomposing the one-step perturbation generation into an iterative process, and proposed I-FGSM and PGD attacks. MI-FGSM [25] introduced a momentum term in iterative attacks, which accumulates previous gradient vectors to guide the calculation of the next gradient. Subsequent methods such as NI-FGSM [26] and VMI-FGSM [37] were also developed. Although these methods achieve high attack success rates on white-box models, their transferability to black-box models is often limited due to the significant differences in gradient information between models.
II-B Input Transformation-based Attack Methods
Methods based on input transformations apply certain transformation enhancement operations to the input examples, increasing the diversity of examples fed into the model. This approach identifies model-discriminative regions common within the image, mitigating the overfitting of adversarial examples to the surrogate model. The DI [27] method is one representative approach. RDI [38] extends the DI approach by resizing the expanded image to its original size. TI [39], SIM [26], and Admix attack [28] are classical input transformation techniques that enhance diversity. Dong et al. [40]proposed a superpixel-guided attention attack method, where perturbations are no longer added at the pixel level, thereby improving the robustness of adversarial examples against defenses based on image processing. Long et al. [41] diversified inputs by performing transformations in the frequency domain. Wang et al. [42] addressed the limitations of existing input transformation methods ([26, 28]) in terms of adaptability and feature fusion effectiveness. Unlike previous methodologies, Guo et al. proposed MixCam [43], which enhances the transferability of adversarial examples from the perspective of shifting the regions that the model focuses on. Additionally, Wang et al. introduced BSR [44], improving the consistency of adversarial example attention heatmaps across different models. Liu et al. [45] designed a method with dynamically adjusted adaptive step sizes while enhancing input diversity, thereby further boosting the transferability of adversarial examples. These methods effectively improve attack transferability by increasing the diversity of inputs.
II-C Advanced Loss Functions-based Attack Methods
Zhou et al. proposed Transferable Adversarial Perturbations (TAP) [21], which improve transferability by increasing the feature distance between the original and adversarial images in intermediate layers. Later, Ganeshan et al. introduced the Feature Disruptive Attack (FDA) [30], where the features of every layer in the model are disrupted. Inka et al. [31] modified the features of the original image to align its intermediate-layer representation becomes closer to that of another class. On the other hand, Wu et al. [46] approached the problem by extracting model attention, prioritizing the disruption of key features that may be adopted by multiple model architectures. Li et al. [47] introduced a method that mixes features of clean and adversarial examples, attenuating intermediate-level perturbations from clean features. Additionally, Wang et al. [48] proposed Diversifying the High-level Features to increases the diversity of features. Since the mapping distance of intermediate features correlates with the distance of the final output of neural networks [21], these methods increase the likelihood of erroneous predictions and enhance the success rate of black-box transfer attacks.
Figure 1: The overall workflow of the method. Before feeding the image into the substitute model, carefully crafted adversarial examples are partitioned in the spatial and spectral dimensions. Each block undergoes random transformations to enhance the diversity of the input. Subsequently, both adversarial and benign examples are input into the substitute model to extract feature maps from the intermediate layers. The weighted feature distancing loss and classification loss are computed to obtain the average gradient, which is iteratively updated to refine the perturbation. Here, $F_{x}^{l}$ and $F_{x_{adv}}^{l}$ denote the features of the benign and the adversarial examples at layer $l$, respectively.
II-D Adversarial Attacks and Defense in Remote Sensing
Chen et al. [49] analyzed the vulnerability of CNN-based remote sensing image (RSI) recognition models to adversarial examples and found that models with different architectures exhibit varying degrees of vulnerability when trained on the same RSI dataset. Later, Xu et al. [50] and Chen et al. [51] investigated the performance of classic attack methods in remote sensing scene classification tasks. Additionally, Burnel et al. [52] used a Wasserstein Generative Adversarial Network-based approach to generate natural adversarial examples. Shi et al. [53] studied the vulnerabilities of multi-modal RSI classification networks and proposed a Multi-Feature Collaborative Adversarial Network. Further studies [33] explored the transferability of adversarial examples in HSIs by updating perturbations across spatial and spectral dimensions, these studies produced smoother adversarial examples while enhancing robustness against defense models. Previous studies that typically added perturbations globally, Liu et al. [54] achieved effective local pixel attacks, balancing attack effectiveness and cost. To mitigate the threat of adversarial attacks, various defenses have been proposed. Notably, Shi et al.[55] introduced an attention-based adversarial defense model that extracts and separates attack-invariant and attack-variant features from adversarial examples, using an attention-guided reconstruction loss to address the attention shift caused by perturbations. Xu et al. [56] leverage the global context information of HSIs to significantly improve model robustness against adversarial examples. They also proposed the Spatial-Spectral Self-Attention Learning Network [57], which integrates spatial and spectral information through a multi-scale spatial attention module and global spectral transformer. To the best of our knowledge, previous studies, with the except for [33], have not yet deeply explored how to improve the transferability of adversarial attacks in HSI classification, which we believe is also of significant research importance.
III Proposed Method
III-A Preliminary
In this study, we refer to the known model used to train and generate adversarial examples as the substitute model, denoted as $f$. The input HSIs are denoted as $x\in X\subset{R^{C\times H\times W}}$, where $H$ and $W$ represent the height and width, respectively, and $C$ is the number of spectral bands. The ground-truth label of $x$ is $y$. The operation of transforming data is denoted as $T\left(\cdot\right)$, and the adversarial example is $x_{adv}\in X\subset R^{C\times H\times W}$. The goal of this study is to find a small perturbation $\delta\subset R^{C\times H\times W}$ that satisfies the constraint condition and enables the adversarial example $x_{adv}=x+\delta$ to maximally mislead the model’s predictions. The loss function is denoted as $L\left(\cdot\right)$ (typically using the cross-entropy loss), so the process of generating adversarial examples is naturally achieved by maximizing the loss function, and it can be expressed as:
${x_{adv}}=\arg\max L\left(f\left(T\left(x_{adv}\right)\right),y\right),s.t.{\left\|\delta\right\|_{p}}\leq\varepsilon$ (1)
Typically, the $l_{p}$-norm is used to constrain the perturbation $\delta$ (in our work, the $l_{1}$-norm is utilized).
III-B Overall Process
The pipeline of our method is illustrated in Fig. 1, The transferability of adversarial examples is enhanced through two key steps: random spatial-spectral block transformation and feature distancing loss. Before feeding the perturbed adversarial examples into the substitute model, $n$ copies are created. These $n$ copies are randomly divided along both the spatial and spectral dimensions. Subsequently, each block will randomly apply a data augmentation method to generate a set of augmented transformed images, which will serve as input for the alternative model. To guide perturbation updates, we integrate intermediate feature distance information between the original and adversarial examples with output layer classification loss to formulate the total loss. Specifically, we first enlarge the original features of the original examples and compute the variance of each channel. The variances of the channels are used to calculate channel weights, which are then incorporated into the weighted Chi-squared distance between the original and adversarial features. The auxiliary classification loss is combined with the feature distancing loss as the total loss function. Because the transformed images may differ significantly from the original ones, the calculation of gradients could introduce instability. To address this, we compute the average gradient across multiple transformed images. In this way, during the iterative optimization process, the perturbation will be updated in the direction of maximizing the feature distance and the prediction distance between the original and adversarial examples, thereby enhancing transferability. The details are introduced below. Algorithm 1 outlines the perturbation generation process integrated with the baseline method MI-FGSM [25].
III-C Block Random Transformation
Data augmentation methods can mitigate the issue of adversarial examples overfitting to the substitute model. Therefore, to enhance the transferability of HSIs, input transformation is essential. Unlike some previous methods, we also consider the structural information of the image. Specifically, given an image $x$, we randomly divide it into $a\times a$ blocks, and the relative relationships among these anchors are the structure of the image, where each anchor is the center of a image block. This structure captures important semantic information for human recognition. Under the premise of preserving the image structure, the transformation of image blocks does not impair recognition but instead generates more diverse images, which is beneficial for improving transferability. On hyperspectral datasets, if we do not disrupt the overall relative structure of the image and only perform block-wise transformations in the spatial dimension, we find that the adversarial examples generated in this way improve the attack transfer success rate compared to baseline methods. However, when compared to the spatial attention maps across various models, the spectral attention maps exhibit more significant differences[33]. Different classification models focus on distinct discriminative spectral bands, so considering spectral neighborhood information may improve transferability.
Figure 2: (a) Schematic of image division. The division parameters in both the spatial and spectral dimensions are set to 2, resulting in 2×2×3 blocks; (b) Schematic of image transformation. Taking spatial transformation as an example, the image is divided into 2×2 blocks, with different transformations applied to each block as shown.
Therefore, in order to further improve the diversity of HSIs, we make full use of the redundancy of the spectral bands in HSIs. First, a set of images is obtained by copying the adversarial examples. On the basis of following the idea of structural invariance, each image is randomly partitioned into $a\times a$ blocks in the spatial dimension and $b$ blocks in the spectral dimension, resulting in $a\times a\times b$ blocks (the division diagram is shown in Fig. 2. Then transform it block by block, randomly select different transformation methods to apply to the image block, and finally get a set of transformed images (the transformation diagram is shown in Fig. 2).
Algorithm 1 Ours Adversarial Attack Method
  Input: HSI example $x$, surrogate model $f$, loss function $L$, perturbation budget $\delta$, step size $\alpha$, decay factor $\mu$, momentum term $m_{i}$, gradients $g_{i}$, number of copies $N$, iteration times $I$.
  Output: HSI adversarial example $x^{adv}$.
1:  $\delta_{0}=\text{zero}(C\times H\times W)$, $x_{0}^{adv}=x$, $g_{0}=0$, $m_{0}=0$
2:  for $i=0\to I-1$ do
3:     Construct a set $X^{adv}_{aug}$ of $N$ transformed images using $T()$:
4:            $X^{adv}_{aug}=T(x+\delta_{i})^{*}N$
5:     Calculate the channel weights according to (3);
6:     Calculate the feature distancing loss according to (4);
7:     Calculate the auxiliary loss according to (5);
8:     Then resulting in the total loss:
9:            $L=L_{F}+\eta*L_{A}$
10:     Calculate the average gradient of $N$ copies:
11:           $g_{i}^{ave}=\frac{1}{N}\left(\sum_{j=0}^{N}(\nabla_{\delta}L(x_{j}^{adv},x))\right),(x_{j}^{adv}\in X^{adv}_{aug})$
12:     Update the momentum:
13:           $m_{i+1}=\mu\cdot m_{i}+\frac{g_{i}^{ave}}{||g_{i}^{ave}||_{1}}$
14:     Update perturbations:
15:           $\delta_{i}={clip_{\varepsilon}}\{\delta_{i}+\alpha\cdot{sign}(m_{i+1})\}$
16:     Update adversarial examples: $x_{i+1}^{adv}=\text{clip}_{x}\{x+\delta_{i}\}$
17:  end for
18:  output $x^{adv}=x_{I}^{adv}$
TABLE I: The various transformation methods adopted in our approach and the resulting transformed images
Raw VShift VFlip Scale Rotate Resize HShift HFlip Dropout DCT AddNoise
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
The methods used are summarized in Table I. In order to avoid information loss, we add constraints on some transformations, such as limiting the scope of scale transformations, preventing over-dark images. By transforming blocks from different spectral bands within the same spatial region, we increase the diversity of transformations is increased. This forces the model to focus on a broader range of spectral information for correct classification, thereby enhancing the impact of perturbations in the spectral dimension and improving the effectiveness of adversarial examples.
(a) raw
(b) channel 10 (0.1213)
(c) channel 28 (0.2106)
(d) channel 50 (0.3260)
Figure 3: Feature maps of different channels of the image within the same layer of the same model, along with their variances.
III-D Feature Distancing Loss
Current research on adversarial examples for RSI classification rarely considers the image features extracted by the model. However, these features are transferable between models with different architectures [29]. Relying exclusively on information from the classification-layer is inadequate for effectively disrupting the deep-level features of the examples. Therefore, we believe that leveraging more feature information during the attack phase could improve adversarial transferability.
Based on this, we introduce the distance between the feature mappings of the original and adversarial examples as the primary guiding loss. This way, the gradient of the intermediate loss with respect to $x$ can maximize the loss in fewer steps, thereby mitigating the vanishing gradient problem. The goal at this stage is to increase the distance between the features of the original and adversarial examples, thereby minimizing the probability that the feature distribution of the adversarial example belongs to a specific class. As the intermediate feature mapping distance between $x$ and $x_{adv}$ increases, it also leads to a larger distance in the final output of the neural network, thereby increasing the loss function and causing mispredictions [21]. We refer to the process of maximizing feature loss as the distancing operation between image features.
III-D1 Weighted Feature Distancing
When using the feature distancing loss, we can choose to make the features of all layers distant, just like TAP. However, due to differences in model architectures, the early layers of some models fail to learn truly salient features and semantic concepts, while the later layers exhibit strong model-specific features[29]. In contrast, the intermediate layers tend to possess well-separated class representations and are less correlated with the model’s architecture. Therefore, we compute the feature distance of the intermediate layers as the loss. Even when extracting features from the same convolutional layer of a model, significant differences can still be observed between the feature maps of different channels, and the variance of features in each channel also varies (variance is a statistic reflecting the richness of texture information in an image or feature), as shown in Fig. 3. When the variance in a specific channel is large, the paper [36] suggests that the features extracted by that channel are non-robust features related to the true class, which have a greater impact on the model’s classification results.
If the distance is calculated by simply averaging the features across all channels, this variability is overlooked. Therefore, to make the perturbation more effectively disrupt the features within the adversarial example, we calculate the weight of each channel based on the magnitude of its variance, thereby giving non-robust features a greater influence in the loss. This may increase the probability of the adversarial examples deviating from the ground-truth label. Let the intermediate layer $l$ of the substitute model $f$ be denoted as $f^{l}$. Assume $A^{l}\in R^{C_{l}*D_{l}}$represents the features of the adversarial example $x_{adv}$ at layer $l$, where $C_{l}$ is the number of channels and $D_{l}$ is the product of width and height. Similarly, $X^{l}\in{R^{C_{l}*D_{l}}}$ represents the features of the original clean example $x$ at layer $l$. Incorporating the feature weights defined by the variance ratio into the distance function. Let $\sigma_{k}^{X^{l}}$ denote the variance of the $k$-th channel in $X^{l}$, which can be expressed by the following formula:
$\sigma_{k}^{X^{l}}=\frac{1}{D_{l}}{\sum\limits_{i=1}^{D_{l}}{\left({X_{k,i}^{l}-\mu_{k}^{X^{l}}}\right)}^{2}}$ (2)
$\mu_{k}^{X^{l}}=\frac{1}{D_{l}}\sum\nolimits_{i=1}^{{D_{l}}}{X_{k,i}^{l}}$ , $X_{k,i}^{l}$ denotes the $i$-th feature value in the $k$-th channel of $X_{l}$.
Then, we can calculate the weight for each channel, which is the proportion of the variance of a single channel to the total variance across all channels. This can be expressed as follows:
${W_{k}}=\frac{\sigma_{k}^{X^{l}}}{{\sum\nolimits_{k=1}^{C_{l}}{\sigma_{k}^{X^{l}}}}}$ (3)
When $W_{k}$ is higher, it shows that the model pays more attention to the $k$-th channel.
For each convolutional layer in a CNN, convolutional kernels extract various features from the input. There is an overall correlation among features across different channels, such as feature distribution, but Euclidean distance ignores such correlations. Furthermore, the Chi-squared distance provides a fair starting point for evaluating the significance of differences between each pair of channels. Therefore, we introduce weights into the Chi-squared distance to compute the feature difference between $x$ and $x_{adv}$. It can be expressed as:
$L_{F}\left({x_{adv},x}\right)$ $=W\cdot d\left(f^{l}\left(x_{adv}\right),f^{l}\left(x\right)\right)$
$=W\cdot d({A^{l}},{X^{l}})$ (4)
$=\sum\limits_{k=1}^{C_{l}}{W_{k}}\frac{M_{k}}{N_{k}}$
where
[…]
We define the above loss as the objective for generating adversarial examples. However, due to the pooling layers in CNNs, some information is inevitably lost, and certain intermediate layer features have lower correlation with the true class of the image [58]. This limitation reduces the attack success rate when relying solely on intermediate layer information. To address this, we integrate output layer information into an auxiliary loss:
${L_{A}}\left(x_{adv},x\right)=CE\left(f\left(x_{adv}\right),f(x)\right)$ (5)
Thus, by combining the previous $L_{F}$ with the auxiliary loss, we obtain the total loss, which can be expressed as follows:
$L={L_{F}}+\eta*{L_{A}}$ (6)
where $\eta$ is a hyperparameter representing the proportion of the auxiliary loss.
By using this loss to iteratively update the perturbation, this method makes more comprehensively utilizes the feature information extracted by the model, causing the feature maps of the adversarial examples to move away from the high confidence regions of the true class. Additionally, since different channel features have varying impacts on the classifier’s prediction results, it achieves more effective adversarial perturbations.
III-D2 Enlargement of the original examples features
If we directly increase the feature distance between the original and adversarial examples, the feature values of the adversarial examples at corresponding positions may either increase or decrease. This imbalance in variation could lead to oscillation issues during gradient updates, preventing effective disruption of features related to the true class. Therefore, we employ a feature enlargement method, scaling up the feature values of the original examples by a certain factor to make them larger than those of the initial adversarial examples. In this way, when optimizing based on the target function to achieve the effect of feature distancing, the feature values at the corresponding positions of the adversarial examples should be updated in the direction of decrease. Although transformation operations are applied to adversarial examples, Ding et al. [59] pointed out that data augmentation operations only alter the spatial positions and areas of different feature components in the feature map without changing their numerical relationships. Thus, overlapping regions can still suppress the feature values within adversarial examples. By averaging the gradients of multiple adversarial examples, the perturbation more stably suppresses the features related to the ground-truth label in these adversarial examples, thereby further enhancing the effectiveness of the feature distancing strategy. The updated formula is as follows:
${L_{F}}\left(x_{adv},x\right)$ $=W\cdot d\left({f^{l}}\left(x_{adv}\right),\lambda*{f^{l}}\left(x\right)\right)$ (7)
$=W\cdot d\left({A^{l}},\lambda*{X^{l}}\right)$
where $\lambda$ is the scaling factor for enlarging the feature values of the original examples. Thus, the optimization formula:
$x_{adv}=\arg\max\left({L_{F}}+\eta*{L_{A}}\right),s.t.\left\|\delta\right\|\leq\varepsilon$ (8)
This method not only preserves the diversity introduced by data augmentation but also enhances the destructive capability of adversarial examples.
IV Experiments
In this section, we detail the implementation details and experimental results of our method on two public datasets. First, we introduce the dataset information, parameter settings, and comparison methods. Next, we evaluate the transferability of the adversarial examples on black-box models and their robustness against two defense methods. Finally, we analyze the impact of several hyperparameters on the experimental results.
IV-A Experimental Setup
1) Datasets: Experiments were conducted on two widely recognized public datasets: HoustonU 2018 and PaviaU, as shown in Fig. 4.
Refer to captionRefer to captionRefer to caption
Figure 4: False color images and ground-truth maps of the (a) HoustonU 2018 and (b) PaviaU datasets.
TABLE II: Training and testing sets of the HoustonU 2018
and PaviaU datasets
HoustonU 2018 PaviaU
Class Train Test Class Train Test Class Train Test
1 5000 4799 11 1000 24002 1 900 5731
2 10000 22502 12 100 516 2 900 17749
3 500 184 13 10000 36358 3 900 1199
4 10000 3588 14 5000 4849 4 900 164
5 3000 2048 15 5000 1937 5 900 445
6 3000 1516 16 10000 1475 6 900 4129
7 200 66 17 100 49 7 900 430
8 10000 29762 18 3000 3578 8 900 2782
9 10000 213684 19 3000 2365 9 900 47
10 10000 35810 20 3000 3824
HoustonU 2018: This dataset was captured by the HSI CASI 1500 sensor in the Houston University area on February 16, 2017. It has a spatial resolution of 1 meter, a spatial size of 601 × 2384 pixels, 48 spectral bands, and 20 available classes of interest. The dataset can be obtained by applying through the official website.
PaviaU: This dataset was collected by the ROSIS-03 satellite over the Pavia University area. It has a spatial resolution of 1.3 meters, a spatial size of 610 × 304 pixels, and includes 115 spectral bands. Prior to the experiments, 12 noisy spectral bands were removed, leaving 103 usable spectral bands and 9 available classes for classification.
2) Training examples: For each pixel, a neighborhood window with a spatial size of 32 × 32 is applied to construct the HSIs examples. The detailed number of training and testing examples for both datasets is provided in Table II.
3) Networks: We use five pretrained classification networks to verify the transferability of the attack: VGG-11, Resnet-18, VGG-19, Inception v3 (Inc-v3), and Inception Resnet V2 (IncRes-v2). Consistent with the paper [33], VGG-11 and Resnet-18 served as white-box models to generate adversarial examples, and other models are used to test transferability. In addition, we also choose two common defense methods based on image processing, which are random noise addition and spectral smoothing filtering, to test the robustness of the attack.
4) Attack Methods: We use FGSM and MI-FGSM as baseline attack methods, and combine our method with baseline attack methods, namely ours-FGSM and ours-MI-FGSM. Additionally, the study [33] addresses related issues and is denoted as SS-FGSM and SS-MI-FGSM. The most relevant work [35] is denoted as SIA-FGSM and SIA-MI-FGSM, respectively. Our method is compared against these approaches to demonstrate its effectiveness.
TABLE III: Intermediate layers used for feature extraction. ”Block” refers to the basic blocks in ResNet-18. ”ReLU-k” denotes the k-th ReLU layer in VGG-11
Model Layer1 Layer2 Layer3 Layer4
Resnet-18 Block-2 Block-4 Block-6 Block-8
Vgg-11 ReLU-1 ReLU-2 ReLU-4 ReLU-6
TABLE IV: Classification accuracy (%) of adversarial attacks on five trained networks on the HoustonU 2018 dataset. The adversarial examples are crafted on ResNet-18. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack Resnet-18 VGG-11 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budge
NONE OA 98.42 98.42 98.67 98.67 98.66 98.66 98.61 98.61 98.63 98.63
AA 98.52 98.52 98.35 98.35 98.61 98.61 98.62 98.62 98.81 98.81
Kappa 97.67 97.67 98.04 98.04 98.02 98.02 97.94 97.94 97.97 97.97
FGSM OA 29.70 19.19 94.71 61.09 94.16 62.60 94.50 61.05 90.99 56.36
AA 16.24 4.83 92.91 28.62 94.27 30.46 89.69 27.03 82.39 20.54
Kappa 8.79 -0.72 92.20 37.83 91.38 41.02 91.86 35.38 86.75 33.92
SS-FGSM OA 15.90 4.20 95.24 69.58 94.06 67.85 94.98 70.72 91.33 60.96
AA 5.22 0.43 94.32 50.71 94.48 57.30 91.29 45.44 86.31 34.86
Kappa -3.10 -10.90 92.98 55.38 91.25 53.10 92.58 55.89 87.28 44.72
SIA-FGSM OA 17.60 2.71 89.61 49.37 88.57 49.00 88.39 55.64 81.80 42.79
AA 6.42 0.43 86.11 27.91 90.27 35.36 80.49 27.79 68.27 20.00
Kappa -3.22 -11.93 84.55 28.26 83.25 28.80 82.55 31.71 72.96 20.40
ours-FGSM OA 26.61 7.58 88.89 44.67 88.44 44.47 86.40 53.07 80.23 40.11
AA 9.30 0.70 81.80 21.02 90.31 27.83 72.22 20.74 62.49 16.85
Kappa -0.05 -13.04 83.35 19.67 83.11 21.05 79.17 24.41 70.28 14.59
MI-FGSM OA 10.60 1.75 94.06 57.16 92.63 54.04 93.89 59.96 89.02 48.37
AA 3.16 0.56 93.04 37.11 93.40 43.57 89.93 33.58 81.88 26.67
Kappa -5.48 -10.94 91.24 38.65 89.17 35.26 90.97 39.50 83.89 30.05
SS-MI-FGSM OA 13.15 4.50 93.34 53.35 91.82 53.42 92.89 57.84 87.77 47.09
AA 3.96 0.57 91.39 30.77 92.76 34.16 86.63 31.83 77.15 23.41
Kappa -4.40 -10.40 90.17 31.90 87.95 32.69 89.46 34.76 82.03 25.79
SIA-MI-FGSM OA 15.49 2.73 85.99 37.46 84.93 35.17 85.01 42.54 77.50 32.46
AA 5.52 0.71 79.47 15.09 87.30 16.57 73.84 16.15 61.52 12.13
Kappa -4.80 -12.58 79.23 13.43 78.00 12.00 77.45 14.74 68.85 8.96
ours-MI-FGSM OA 25.15 9.11 81.73 33.58 82.22 31.93 79.95 41.36 72.17 30.58
AA 7.80 0.84 70.62 11.32 85.83 11.73 62.74 12.20 51.63 10.10
Kappa -2.00 -13.73 72.84 5.04 74.68 4.60 69.30 8.89 58.66 2.81
5) Parameters: In this paper, two smaller disturbance budgets are selected for experiments, which are 0.01 and 0.03 respectively. Some parameters of the SS-FGSM method follow the settings of the original paper: the spatial superpixel size is set to four spatial pixels, and the spectral clustering size is set to 20. The ground sample distances (GSDs) of the HoustonU 2018 dataset is 4 m, and that of the PaviaU dataset is 5.2 m. The number of external iterations $T=20$, $\mu=1$. Considering that the improvement in accuracy is not evident relative to the computational complexity, the parameter $t$ is set to 1. To ensure fairness with other methods, the main parameters of our method are established as follows: the spatial division parameter is set to 3, consistent with SIA, and the spectral division parameter is also set to 3. The iteration count is set to 20, the number of transformed example copies is 10 and the feature loss weight coefficient $n$ is set to 0.1. For the white-box model, one layer is selected for feature extraction from shallow to deep layers (1/2/3/4, as shown in Table III). In our experiments, both models selected layer 3. The original feature enlargement factor is set to 1.2, the momentum factor for MI is set to 1, the step size is 2/255, and the batch size is 100. The experiments are carried out with 13th Gen Intel Core i5-13600KF CPUs with 32 GB of RAM, and one NVIDIA GeForce RTX 4090 GPU.
TABLE V: Classification accuracy (%) of adversarial attacks on five trained networks on the HoustonU 2018 dataset. The adversarial examples are crafted on VGG-11. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budge
NONE OA 98.67 98.67 98.42 98.42 98.66 98.66 98.61 98.61 98.63 98.63
AA 98.35 98.35 98.52 98.52 98.61 98.61 98.62 98.62 98.81 98.81
Kappa 98.04 98.04 97.67 97.67 98.02 98.02 97.94 97.94 97.97 97.97
FGSM OA 66.60 59.61 92.25 69.17 88.71 66.96 92.57 67.54 91.07 64.73
AA 35.66 16.57 88.68 39.13 85.41 30.00 86.18 29.50 81.14 24.14
Kappa 52.25 39.71 88.71 54.55 83.39 49.37 89.06 48.83 86.91 45.97
SS-FGSM OA 53.90 39.83 92.85 68.72 88.22 60.55 92.43 69.96 90.48 67.30
AA 21.30 9.57 82.27 41.70 85.88 36.20 81.33 40.96 73.92 36.21
Kappa 33.76 15.96 89.44 53.26 82.52 41.34 88.69 54.03 85.77 51.18
SIA-FGSM OA 28.04 6.70 80.48 40.50 70.34 33.51 75.41 46.29 70.48 39.51
AA 10.30 0.70 63.98 21.11 60.78 16.88 54.22 21.36 49.54 15.12
Kappa 2.60 -16.08 71.23 14.67 56.93 6.65 62.87 18.85 56.32 11.81
ours-FGSM OA 40.25 15.20 79.94 31.27 73.44 26.71 72.51 37.82 68.43 31.82
AA 19.32 1.40 61.78 12.86 64.47 9.62 50.83 13.35 44.34 9.93
Kappa 9.86 -20.08 69.76 -4.37 60.62 -10.52 56.90 -0.14 51.35 -5.80
MI-FGSM OA 51.96 48.22 91.40 63.98 84.86 56.20 90.96 64.42 88.79 62.69
AA 16.15 10.43 80.02 35.41 80.92 25.55 78.50 31.52 71.25 27.21
Kappa 30.36 25.16 87.32 46.86 77.48 35.71 86.50 46.78 83.27 45.18
SS-MI-FGSM OA 52.92 44.61 90.02 61.79 82.91 55.76 89.35 63.97 87.31 61.58
AA 19.23 10.67 76.15 31.68 76.35 25.26 74.12 32.38 69.36 28.17
Kappa 32.09 21.97 85.29 43.50 74.57 35.15 84.07 45.82 81.09 43.05
SIA-MI-FGSM OA 22.83 6.60 76.03 24.30 64.66 19.92 71.91 26.67 66.09 24.31
AA 8.33 0.68 59.83 11.16 53.93 7.38 50.80 9.64 46.04 6.71
Kappa -1.72 -14.37 65.14 -0.61 49.19 -3.99 58.12 0.68 50.53 -1.82
ours-MI-FGSM OA 32.04 16.21 72.21 23.54 64.83 21.88 66.17 28.55 60.77 25.29
AA 11.80 1.49 54.01 9.30 52.61 6.57 43.54 9.64 38.67 6.67
Kappa -1.09 -18.60 58.48 -10.45 48.11 -13.94 47.49 -8.09 40.57 -10.99
TABLE VI: Classification accuracy (%) of adversarial attacks on five trained networks on the PaviaU dataset. The adversarial examples are crafted on ResNet-18. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack Resnet-18 VGG-11 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budge
NONE OA 99.96 99.96 99.85 99.85 99.95 99.95 99.81 99.81 99.94 99.94
AA 99.97 99.97 99.88 99.88 99.96 99.96 99.81 99.81 99.92 99.92
Kappa 99.94 99.94 99.84 99.84 99.92 99.92 99.80 99.80 99.91 99.91
FGSM OA 39.11 12.49 78.77 35.22 80.91 56.91 81.86 53.48 95.74 62.70
AA 58.24 31.82 89.91 51.05 93.88 64.57 94.86 64.56 98.83 82.32
Kappa 27.87 1.14 70.97 20.34 74.25 44.18 75.73 42.67 93.92 53.69
SS-FGSM OA 20.67 2.18 51.83 18.74 62.30 33.55 75.92 48.05 89.06 52.22
AA 43.09 15.62 82.37 43.70 84.41 55.66 93.70 62.61 97.51 76.86
Kappa 6.39 -14.26 40.51 3.96 52.91 19.74 68.93 36.93 84.96 43.12
SIA-FGSM OA 16.92 0.72 44.02 15.91 57.20 35.87 73.08 43.05 88.27 48.24
AA 39.40 7.05 75.34 42.25 84.63 62.72 91.17 60.98 97.23 73.37
Kappa 1.09 -11.98 30.87 1.96 46.23 24.24 65.13 31.60 83.93 38.30
ours-FGSM OA 20.83 1.84 41.04 9.44 55.43 29.53 72.06 40.55 88.82 46.26
AA 43.96 12.11 72.77 32.16 83.28 50.42 90.25 57.14 97.27 72.30
Kappa 5.35 -11.31 27.58 -6.73 43.66 16.47 63.53 28.14 84.62 36.12
MI-FGSM OA 20.48 1.83 55.87 24.54 65.41 40.76 75.60 46.88 89.25 52.18
AA 41.91 13.70 83.56 47.76 90.12 57.82 93.35 61.96 97.46 75.71
Kappa 7.89 -10.20 45.73 11.91 56.98 28.49 68.59 36.93 85.20 43.11
SS-MI-FGSM OA 22.63 8.04 54.06 17.85 64.55 33.97 76.19 45.80 89.13 50.98
AA 44.99 26.41 84.52 42.50 90.47 53.85 93.93 61.42 97.52 73.17
Kappa 9.40 -6.46 43.67 4.47 56.16 21.19 69.29 35.37 85.05 41.90
SIA-MI-FGSM OA 18.16 0.84 43.95 12.10 56.06 30.55 72.28 40.48 88.23 48.06
AA 40.21 7.51 75.30 35.89 83.49 49.09 90.37 57.54 97.13 71.91
Kappa 2.66 -12.14 30.77 -3.84 44.47 15.64 63.98 28.96 83.86 38.54
ours-MI-FGSM OA 20.52 2.08 40.47 6.18 54.40 25.69 71.08 37.06 88.57 45.90
AA 43.75 12.55 72.62 28.51 82.44 45.06 89.32 54.23 97.11 71.95
Kappa 5.16 -12.23 27.09 -11.27 42.23 9.89 62.10 23.81 84.28 35.83
IV-B Transferability Experiments
In this section, we evaluate the performance of the proposed method on five pretrained classification models. The evaluation metrics include Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient. Lower values of these metrics indicate a lower recognition accuracy of the model on adversarial examples, thereby reflecting better adversarial attack performance. The experimental results on the HoustonU 2018 dataset are presented in Tables IV and V, while the results on the PaviaU dataset are shown in Tables VI and VII. The maximum perturbation budgets used were 0.01 and 0.03. The experimental results represent the averages of multiple trials to reduce the influence of randomness introduced by random operations.
HoustonU 2018 Dataset: The experimental results are shown in Tables IV and V. Adversarial examples were generated using the white-box models ResNet-18 and VGG-11, respectively. It can be seen that across all methods and models, the attack effect of the perturbation budget of 0.03 is significantly better than the perturbation budget of 0.01, indicating that smaller perturbation budget pose greater challenges. When combined with MI-FGSM, our method achieved comparable or even better improvements at a budget of 0.01 compared to 0.03. This indicates that our method can effectively enhance attack performance even under smaller perturbation budgets. The attack performance of our method across all models is significantly better than that of the baseline method. For a perturbation budget of 0.01, the average improvement was 14.06%, while for a budget of 0.03, the average improvement was 25.5%. Compared to the SIA method on four black-box models, the attack success rate increased by an average of 2.44% at $\delta=0.01$ and 3.28% at $\delta=0.03$. However, we observed that the white-box attack success rate decreased significantly compared to SIA. Data from ablation experiments indicate that this is caused by spectral-wise block transformations and feature distancing loss. We believe that the possible reason is that the use of multi-dimensional block transformations and weighted Chi-squared distance loss alter the adversarial example optimization process. This approach not only focuses on the classification boundary of the surrogate model but also pay more attention to the change of feature distribution. Consequently, this may lead to a more dispersed disturbance in the input space, thereby diminishing the effectiveness of attacks on the substitute model and resulting in a poorer performance in white-box attacks.
TABLE VII: Classification accuracy (%) of adversarial attacks on five trained networks on the PaviaU dataset. The adversarial examples are crafted on VGG-11. The maximum perturbation budget is set to 0.01 and 0.03, respectively
Attack VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
Method Perturbation 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03 0.01 0.03
budge
NONE OA 99.85 99.85 99.96 99.96 99.95 99.95 99.81 99.81 99.94 99.94
AA 99.88 99.88 99.97 99.97 99.96 99.96 99.81 99.81 99.92 99.92
Kappa 99.84 99.84 99.94 99.94 99.92 99.92 99.80 99.80 99.91 99.91
FGSM OA 32.18 12.14 70.14 37.18 66.34 50.74 83.09 60.64 93.62 55.37
AA 46.52 28.52 86.92 63.42 86.27 66.21 94.89 74.77 98.46 77.26
Kappa 14.02 -4.08 60.05 24.10 55.30 37.76 77.05 49.53 91.00 45.98
SS-FGSM OA 19.18 8.72 60.23 34.14 60.50 40.97 80.75 62.27 91.41 57.49
AA 37.98 17.83 83.82 62.08 86.04 65.64 95.51 78.27 98.03 80.21
Kappa 2.08 -8.65 49.20 20.98 49.60 27.48 74.59 53.01 88.04 49.33
SIA-FGSM OA 13.22 0.79 46.69 20.92 52.92 32.26 73.51 49.01 87.22 53.06
AA 32.77 6.06 69.30 42.60 80.84 50.55 90.93 64.39 96.96 71.35
Kappa -3.44 -16.61 33.82 4.28 40.04 16.13 65.34 36.71 82.59 43.97
ours-FGSM OA 13.47 0.04 41.82 11.21 49.93 18.46 71.61 38.51 86.95 43.16
AA 35.82 1.49 66.05 35.73 80.24 38.41 88.67 55.40 96.82 63.88
Kappa -2.89 -19.47 28.90 -4.93 37.06 0.88 62.78 23.53 82.23 31.86
MI-FGSM OA 11.56 9.70 61.46 38.16 61.18 42.46 79.79 59.40 87.75 55.70
AA 27.09 12.77 82.00 60.71 85.29 63.10 95.21 75.01 97.19 79.02
Kappa -4.85 -7.65 50.48 23.58 50.05 27.60 73.42 46.81 83.27 47.53
SS-MI-FGSM OA 15.62 9.25 57.43 25.34 59.06 31.70 78.86 50.26 89.30 50.47
AA 35.29 22.38 82.42 52.71 85.32 57.06 94.88 68.43 97.55 73.40
Kappa -1.07 -7.83 46.22 11.20 47.91 16.83 72.29 38.59 85.27 41.18
SIA-MI-FGSM OA 10.35 0.86 44.37 19.62 51.85 24.02 71.67 41.19 85.59 47.46
AA 27.02 6.87 66.51 38.58 80.00 42.29 88.07 57.06 96.46 65.00
Kappa -6.81 -16.26 31.06 2.34 38.78 7.27 62.88 27.51 80.51 37.49
ours-MI-FGSM OA 8.98 0.04 37.40 9.01 48.65 11.73 70.01 31.90 85.83 37.98
AA 29.32 1.56 61.14 33.61 79.04 31.88 85.12 50.42 96.26 57.57
Kappa -7.72 -19.16 23.88 -7.38 35.56 -6.22 60.69 16.04 80.77 25.27
PaviaU Dataset: The experimental results are shown in Tables VI and VII. We present the results of adversarial examples generated on the ResNet-18 and VGG-11 models. Unlike the HoustonU 2018 dataset, attacking black-box models on this dataset is more challenging due to its higher number of spectral bands. On this dataset, our method largely retained the attack success rate on white-box models while significantly improving baseline attack transferability. Moreover, the performance improvement on the PaviaU dataset was slightly better than that on the HoustonU 2018 dataset. Compared to the baseline method across all models, the attack success rate of our method increased by an average of 13.12% at $\delta=0.01$ and 18.07% at $\delta=0.03$. Compared to the SIA method across four black-box models, the attack success rate increased by an average of 2.05% at $\delta=0.01$ and 7.31% at $\delta=0.03$. Compared to ResNet-18, the adversarial examples generated when VGG-11 is used as a substitute model show greater improvements in attack performance. We believe that this is due to differences in model structure, feature extraction mechanisms, and attention areas, which lead to varying attack effects.
IV-C Defense Experiments
To verify the robustness of the proposed method against common defense strategies, we conducted experiments on two public datasets using two defense methods: adding random noise and spectral dimension filtering. Both defense techniques are commonly used in image processing to mitigate or disrupt adversarial perturbations, thereby improving model robustness. Comparisons with the baseline, SS, and SIA methods are presented in Tables VIII and IX.
Adding random noise: Before feeding examples into the model, random noise is added to the examples to disrupt the structure of adversarial perturbations, reducing the model’s sensitivity to adversarial examples and thus weakening the effectiveness of adversarial attacks. The budget for the added random noise is set at 0.1, and the results are shown in Table VIII. It can be observed that our proposed method maintains good attack performance even under the random noise defense.
Spectral filtering: HSIs contain redundant spectral bands; therefore, local spectral filtering is also a challenging defense method. We perform average filtering along the spectral dimension to validate the robustness of the attack method. The filter size is set to 7, and the results are presented in Table IX. Our method demonstrates strong adaptability and robustness across different datasets and defense strategies.
TABLE VIII: Overall classification accuracy (%) of HSI adversarial examples under added random noise conditions. The adversarial examples for the two datasets are generated on VGG-11 with a maximum perturbation budget of 0.03, and the intensity of the random noise is set to 0.1
PaviaU VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 13.09 37.85 60.61 57.84 46.20
SS-FGSM 9.23 35.42 65.10 62.35 53.65
SIA-FGSM 0.89 21.79 54.87 50.19 52.15
ours-FGSM 0.05 11.54 45.71 39.78 50.30
MI-FGSM 9.79 39.82 69.73 59.66 48.11
SS-MI-FGSM 9.80 26.47 53.80 49.90 53.14
SIA-MI-FGSM 1.04 20.55 46.31 43.06 49.68
ours-MI-FGSM 0.09 9.21 34.55 32.82 45.16
HoustonU 2018 VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 59.96 69.17 67.32 67.36 63.14
SS-FGSM 40.31 69.02 61.85 70.40 66.35
SIA-FGSM 7.15 40.63 33.75 45.64 37.29
ours-FGSM 15.75 32.02 26.79 37.79 30.14
MI-FGSM 48.38 65.16 57.25 65.74 63.51
SS-MI-FGSM 45.26 63.06 56.78 65.07 61.74
SIA-MI-FGSM 7.19 25.10 20.17 26.87 23.60
ours-MI-FGSM 17.13 24.22 21.96 28.60 24.23
IV-D Ablation experiment
In this section, we explore the influence of each component of the generation method on attack performance through ablation experiments. Additionally, we investigate how various hyperparameters affect the experimental results.
TABLE IX: Overall classification accuracy (%) of HSI adversarial examples under spectral filtering conditions. The adversarial examples for the two datasets are generated on VGG-11 with a maximum perturbation budget of 0.03, and the filter size range is set to 7
PaviaU VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 14.01 38.85 47.46 64.86 65.03
SS-FGSM 9.43 35.78 44.46 64.88 64.31
SIA-FGSM 3.20 22.52 40.94 57.32 62.36
ours-FGSM 1.10 12.27 29.29 49.33 54.24
MI-FGSM 10.84 40.79 47.31 65.24 66.78
SS-MI-FGSM 9.85 27.54 36.26 54.27 54.42
SIA-MI-FGSM 1.95 20.59 36.04 51.47 57.75
ours-MI-FGSM 1.09 9.42 21.13 42.96 48.66
HoustonU 2018 VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
FGSM 60.21 70.78 67.79 68.71 65.91
SS-FGSM 40.11 70.59 62.84 71.39 68.28
SIA-FGSM 7.46 42.23 36.68 48.91 41.94
ours-FGSM 15.64 32.81 28.67 39.22 33.28
MI-FGSM 48.29 64.97 57.60 65.78 63.91
SS-MI-FGSM 44.86 62.80 56.90 64.85 62.56
SIA-MI-FGSM 7.40 25.15 21.37 28.27 26.26
ours-MI-FGSM 16.77 24.38 22.81 29.71 26.28
First, we analyze the impact of each component on the attack effectiveness. The experimental results are summarized in Table X. The relevant parameters have been described in the previous section on parameter settings. Although performance slightly varies across different models, the attack effectiveness generally improves with the application of our method. The results show that, compared to the spatial dimension, performing block transformations in the spectral dimension can also reduce the model’s classification accuracy. Combining both approaches improved the transferability on the two datasets by approximately 4% and 6%, respectively. Additionally, integrating feature distancing loss further enhanced the attack transfer success rate of adversarial examples. The performance on the PaviaU dataset was better than on the HoustonU 2018 dataset, which may be related to the data distribution and the size of the regions on which the model focuses. Finally, using the enlarged original example features when calculating the Chi-squared distance can further improve the attack success rate, thereby validating the effectiveness of the method.
TABLE X: Ablation study of the proposed approach. The adversarial examples are crafted on VGG-11, with the integrated baseline method being FGSM. The perturbation budget is set to 0.03. SpatialT represents spatial transformation; SpectralT represents spectral transformation; FDL represents feature distancing loss; and Enlarge represents the operation of enlarging the features of the original sample
Constrain condition PaviaU
SpatialT SpectralT FDL Enlarge VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
0.75 21.03 32.25 48.94 52.89
0.55 16.40 23.66 45.62 48.26
0.41 17.18 25.51 45.06 49.08
0.51 15.54 26.35 43.73 46.64
0.19 12.72 20.20 39.77 43.90
0.04 11.17 18.56 38.40 43.02
Constrain condition HoustonU 2018
SpatialT SpectralT FDL Enlarge VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
6.72 40.51 33.53 46.28 39.65
15.61 36.86 30.26 41.44 36.14
13.41 32.88 28.83 39.28 34.27
16.56 38.94 32.84 44.48 36.69
16.56 32.63 28.16 38.72 33.17
15.18 31.27 26.71 37.79 31.87
To gain insight into the performance of the method, we further analyze the influence of various hyperparameters on the experimental results. It mainly includes parameters such as block division, the number of copies, and the feature enlargement factor. The experiments were conducted mainly on the PaviaU dataset, with VGG-11 selected as the white-box model for generating adversarial examples, while the other models were treated as black-box models. The perturbation budget was set to 0.03.
The impact of the number of iterations on attack performance is shown in Fig. 5. The iterations range from 1 to 60. When the number of iterations exceeds 20 times, the improvement in the attack effect is limited. Therefore, to optimize computational costs while considering the attack’s efficacy, we set the number of iterations to 20.
The impact of the number of copies on attack performance is shown in Fig. 5. The number of replicated copies ranges from 1 to 20. As the number of transformed copies increases, the accuracy decreases most rapidly from 1 to 4, and then gradually stabilizes. This trend indicates that additional replicated examples yield diminishing returns in improving attack performance.
The impact of the feature extraction layer on attack performance is shown in Fig. 5. The fourth layer of the VGG-11 model demonstrates the most effective attack performance, while the first layer of the ResNet-18 model has the best attack effect. This difference is mainly due to the differing attention to feature information across models, which is related to their structure. Therefore, when using this method, the selection of the feature extraction layer should be selected to the specific structure of the surrogate model to achieve better attack performance.
Figure 5: (a) the number of iterations, (b) the number of copies, and (c) the impact of different model feature extraction layers on the attack success rate.
Figure 6: The impact of (a) the weight coefficient of feature loss and (b) the feature enlargement factor of the original examples on the attack success rate across different models.
Figure 7: The impact of using different division parameters in the spatial and spectral dimensions on the attack success rate.
The impact of the weight coefficient of feature loss on attack performance is shown in Fig. 6. A weight coefficient of 0 indicates that no auxiliary information from the output layer is used, which is observed to degrade attack performance. When the weight coefficient is set to 0.03 or 0.05, the attack effect is the best. With 0.07 as the boundary, the weight of auxiliary loss increases and the effect gradually deteriorates. Therefore, setting the weight to 0.03 or 0.05 yields a more effective attack.
The impact of the original feature enlargement factor is shown in Fig. 6, with the enlargement factor ranging from 1 to 3. On the VGG-11 model, when the magnification is 2.7 or 3, the attack performance is the best; however, on the ResNet-18 model, the best performance occurs at a magnification of 1.2. This further shows that the effect of feature loss is affected by the model feature extraction method, indicating that optimal parameters can be selected according to the information of the surrogate model.
The influence of division parameters on attack performance is shown in Fig. 7. The range for spatial division parameters was set to 1-5, while that for spectral division was set to 1-4. As the spatial division number increased to 3, the attack performance improves significantly and then stabilized. Similarly, as the spectral division number increased to 3, the improvement in attack performance gradually diminished. To achieve a balance between computational efficiency and time cost, both the spatial and spectral division parameters were set to 3 in this work.
V Conclusion
For the HSI classification problem, existing methods for generating adversarial examples often suffer from overfitting to the surrogate model and fail to fully utilize the features extracted during training. Building on this, we propose a comprehensive attack method to improve the transferability of adversarial examples in HSI classification tasks. Considering that HSIs have a large number of spectral bands, we divide the input hyperspectral examples into blocks along both spatial and spectral dimensions and apply various transformations to each block. Additionally, we design a weighted feature distancing loss function, which updates the perturbation by calculating the average gradient of multiple transformed images. It effectively suppresses the important features in the adversarial examples and enhances the transferability of adversarial attacks on HSIs. Experiments show that the adversarial examples produced by this method achieve significant transferability on multiple black-box models, which not only deepens the understanding of adversarial vulnerabilities in HSI classification but also provides an important reference for developing more robust defense strategies in this field.
Acknowledgment
Acknowledge for IEEE GRSS IADF as well as the Hyperspectral Image Analysis Laboratory at the University of Houston for providing the HoustonU 2018 dataset. This study was funded by the National Key Laboratory of Integrated Aircraft Control Technology Fund Project (WDZC2024601B02) and the National Natural Science Foundation of China (No.42371433).